A 2005 Honda minivan can travel 240 miles in the city on 4 gallons of gas. Find the amount of gas required for 360 miles of city driving. (hint: use a proportion).
Answer by nerdybill(7384)

A 2005 Honda minivan can travel 240 miles in the city on 4 gallons of gas. Find the amount of gas required for 360 miles of city driving. (hint: use a proportion).
.
Let x = amt of gas (gallons) used to drive 360 miles
then
240/4 = 360/x
cross multiplying:
240x = 4(360)
240x = 1440
x = 6 gallons